OpenStudy (anonymous):

Does the equation represent a direct variation? If so, find the constant of variation. 5x = -3y

OpenStudy (blockcolder):

Since y/x=k, divide both sides by -3x: \[\frac{5x}{-3x}=\frac{-3y}{-3x}\\ \Rightarrow -\frac{5}{3}=\frac{y}{x}=k.\]